# Using LLMs in Generating Design Rationale for Software Architecture Decisions

### v1

@@ 3.3. Dataset Construction
As shown in Fig. 2, we constructed a dataset from SO posts, GitHub issues and discussions to facilitate and evaluate the DR generation for practical architecture decisions. To avoid potential data leakage issues, we constructed a new dataset, as the existing datasets may contain data that was used to train the LLMs used in this study as we discussed in Section 3.2. Note that, in our dataset, all posts, issues and discussions were created after the cut-off dates of the five LLMs used in this study. The dataset used in this study has been provided at (Zhou et al., 2025).
@@ 3.3.1. Data Collection
For data collection, we used a keyword-based method to gather data from SO and GitHub. Specifically, we selected the general terms “architect*” (including “architecture”, “architecting”, and “architectural”) to search posts, issues and discussions. Such selection was also used in a recent study to retrieve architecture related SO posts (de Dieu et al., 2023), and it indicates that these terms are effective in capturing more relevant data compared to other general keywords like “design*”. For SO posts, we utilized an SQL query via Stack Exchange Data Explorer interface (Explorer, nd) to collect posts with at least one answer from SO. For GitHub issues and discussions, we collected issues marked as “closed” with the reason specified as “completed”, and discussions marked as “answered”. Note that, we did not limit our search solely to the tags of SO posts, GitHub issues and discussions, as using tags alone may omit some important data related to software architecture topics (Barua et al., 2014). Therefore, we considered both the body and title of posts, issues and discussions during our data search.
Additionally, we limited the retrieval of the posts, issues and discussions to dates after the knowledge cut-off dates of all LLMs. As presented in Table 1, gemini-1.0-pro has the latest cut-off date among the five LLMs we selected, that is, April 2023. Hence, we restrict the data retrieved from SO and GitHub to content after May 2023. We conducted data collection on June 18, 2024, gathering data from SO and GitHub that contained “architect*” in the content body, tags, or title, starting from May 1, 2023. In total, 692 SO posts, 8,050 issues, and 2,546 discussions were retrieved.
@@ 3.3.2. Data Review and Labeling
  • The posts, issues, and discussions must explicitly include an “architecture-related problem”, and the corresponding “architecture decision” and “design rationale”. This criterion ensures that the subsequent data extraction process can retrieve the necessary information from these data entries.
@@
The formal data labeling was conducted by the first author. Prior to this, the first and fourth authors randomly selected 25 SO posts, 25 GitHub issues, and 25 GitHub discussions for a pilot labeling process to minimize personal bias and ensure adherence to the data labeling criteria. The inter-rater reliability between the two authors was measured by the Cohen’s Kappa coefficient (Cohen, 1960), yielding values of 0.834. This result indicates an almost perfect agreement between the two authors (Landis and Koch, 1977). The second and third authors were then involved to resolve discrepancies between the pilot results of the first and fourth authors and reach a consensus on the four criteria for data labeling. Subsequently, the first author conducted the formal data labeling according to the four criteria. For SO posts, the first author reviewed and labeled all 692 posts, and identified 50 architecture-related posts for our research. Given the large number of collected GitHub issues and discussions, and to maintain a balanced data volume between SO and GitHub sources, we selected the top 25 issues and top 25 discussions that meet the criteria, sorted by “best match”. Consequently, our dataset comprises 100 entries from SO and GitHub. The data labeling results were reviewed multiple times by the first, second, third, and fourth authors. Discrepancies were addressed through discussions to reach a consensus. Specifically, the first author presented the reasons for the inclusion or exclusion of certain data. The four authors then analyzed and discussed the reasons based on the criteria for data labeling, ultimately reaching an agreement. The pilot and formal data labeling results are recorded in our dataset (Zhou et al., 2025).
@@ 3.3.3. Data Extraction
We established a set of data items for data extraction, as presented in Table 2. Data items D1 and D2 are provided as part of the prompt for LLMs and LLM-based agents to generate DR. D3 serves as a reference for evaluation in results analysis, as further elaborated in Section 3.5. The data extraction criteria are as follows:
@@
The first and fourth authors randomly selected 5 SO posts, 5 GitHub issues, and 5 GitHub discussions to conducted a pilot data extraction. In case of any discrepancies, the second and third authors were involved in reaching a consensus. The results show that the three data items in Table 2 can be extracted from our dataset. Subsequently, the first author conducted the formal data extraction from the labeled dataset. Each extracted data item was reviewed multiple times by the first, second, third, and fourth authors to ensure accuracy. Any inconsistencies between the four authors were discussed according to the criteria of data extraction to achieve a consensus. The data extraction results are recorded in our dataset (Zhou et al., 2025).
@@
D1 Architecture Problem The key point(s) of the architecture problem in SO posts, GitHub issues, or GitHub discussions If you wish to access Console, Router, or Controller via the internet, you must open ports to the public internet and assign a public DNS record. This opens Ziti services to DDoS.
D2 Architecture Decision The key point(s) of the architecture decision in SO posts, GitHub issues, or GitHub discussions I propose implementing UDP hole-punching on the Controller and Router and Tunneler so they can see each other without opening the Controller and Router ports to the Internet. And since the Tunneler can now see the Controller and Router, you can also generate JWT locally on the Tunneller/Client.
D3 Design Rationale The key point(s) of design rationale that consists of the arguments for the architecture decision in SO posts, GitHub issues, or GitHub discussions UDP hole-punching in products such as ZeroTier and Nebula that may be useful for OpenZiti: Argument_1: The supernode is a separate stand-alone service that does not influence sensitive services, such as the controller/router that does the main work and handles and enforces zero-trust, routing, and handle management. So, if the supernode is hacked, the other components will not be compromised. Argument_2: There are many supernodes across the Internet, so it is hard to DDoS all of them simultaneously. Argument_3: Even if all supernodes are DDoSed and down, the peers that already established connection do not care about the supernode and continue working without an issue. Argument_4: Only one port is opened, and you must be concerned about being open on the public Internet to one service that is not viral and acceptable if not working for hours or days.
@@ 3.4. Design Rationale Generation
To answer the three RQs (see Section 3.1), we designed prompts for LLMs and LLM-based agents to generate DR based on the data items defined in Table 2. The DR generated with three prompting strategies is recorded in the dataset (Zhou et al., 2025), which are further discussed in Section 4.
@@ 3.4.2. CoT (RQ2)
CoT allows LLMs to progressively demonstrate their thought process before arriving at a final answer, enhancing their reasoning ability and the transparency of their decisions (Wei et al., 2022b). Therefore, in addition to “Architecture Problem” and “Architecture Decision”, we also provided LLMs with a framework for analyzing the rationale in prompts. We designed several versions of the framework for LLMs to analyze the rationale and conducted pilot experiments on 10 randomly selected samples from our dataset. According to the results of the pilot experiments, we identified the most efficient CoT framework with four steps shown in Fig. 4. In Step 1, we ask the LLMs to understand architecture decisions based on the context of architecture problems. In Step 2 and Step 3, we ask the LLMs to analyze the advantages and disadvantages of the architecture decisions, respectively. Finally, in Step 4, we ask the LLMs to present the DR based on the trade-off of the advantages and disadvantages.
It is important to note that we did not provide any other examples in prompts for LLMs. This is because the main purpose of providing additional examples in CoT is to help the model understand the required reasoning process (Wei et al., 2022b). However, the results of the pilot experiments demonstrate that LLMs can generate DR step by step according to the designed CoT framework without any examples, which was further confirmed in the formal experiments. In addition, introducing examples in prompts significantly increases the length of input text, which raises the difficulty for LLMs to understand architecture problems and leads to their generated DR being potentially influenced by specific examples.
@@ 3.4.3. LLM-based Agents (RQ3)
To answer RQ3 in Section 3.1, we designed five LLM-based agents (i.e., Aspect_Identifier, Information_Collector, Aspect_Analyst, Aspect_Reviewer, and Trade-off_Analyst) by utilizing the MetaGPT framework (Hong et al., 2023). MetaGPT utilizes an assembly line approach to assign different roles to various agents, effectively breaking down complex tasks into subtasks that require collaboration among multiple agents. According to the study of Liu et al. (Liu et al., 2024a), the design pattern of MetaGPT is role-based cooperation, which has been employed in many SE studies (e.g., (Jin et al., 2024; Ratnadira et al., 2024)). Currently, no research has reached a widespread consensus on the role delineation for generating DR. Thus, we approached a process of generating DR from the perspective of architects (Jansen et al., 2008; Shahbazian et al., 2018), breaking it down into a set of subtasks assigned to five distinct agents.
@@
  • Aspect_Identifier agent is responsible for identifying the most relevant and important aspects related to certain architecture decisions. Specifically, Aspect_Identifier agent receives the architecture problem and the corresponding architecture decision, and then provides up to six of the most relevant aspects, along with their definitions. We set this upper limit at six aspects according to the results of the pilot experiments on 10 samples from our dataset. The results indicated that without such a restriction, LLMs tend to enumerate too many aspects, which can make it difficult to maintain a clear focus in the subsequent analysis of DR. Additionally, this excessive enumeration significantly increases the costs for LLM-based agents. As a result, we found that limiting the aspects to six is considered a reasonable compromise, ensuring both depth of analysis and a concentrated focus. Then, each defined aspect is transmitted to Information_Collector agent.
@@
  • Information_Collector agent is in charge of gathering relevant information through search engines. We initially intended to enhance the information retrieval capabilities of Information_Collector agent by using Retrieval-Augmented Generation (RAG) (Gao et al., 2024) technique with a unified database. However, the diversity of architecture problems prevented us from providing a single database that could adequately address all issues. Consequently, we chose to dynamically retrieve relevant knowledge from search engines based on specific problems. Specifically, Information_Collector agent receives architecture problems, the corresponding architecture decisions, and aspects predefined by Aspect_Identifier agent. Based on each aspect, Information_Collector generates the three most relevant keywords for search engine queries. We chose DuckDuckGo as our search engine as it offers search results without personalized interference and provides a high level of information reliability (Hannak et al., 2013). For each keyword, Information_Collector agent retrieves the top 20 URLs returned by the search engine. Therefore, for each aspect, a total of 60 relevant URLs are collected (i.e., 3 $\times$ 20 URLs). To prevent the agent from using the original SO posts, GitHub issues, or GitHub discussions as reference information (i.e., data leakage issues), we discard any collected background knowledge whose URLs appear in the dataset and replace them with alternative sources. Then, the agent stores the webpage content from these URLs in Facebook AI Similarity Search (FAISS), a library designed for fast nearest-neighbor search and dense vector clustering (Douze et al., 2024). By using the similarity search function of FAISS, Information_Collector agent selects the 10 most relevant pieces of information from all the collected webpages that are most similar to the given architecture problems and architecture decisions. The 10 pieces of information, are then transmitted to Aspect_Analyst agent as background knowledge along with the aspects and their definitions identified by Aspect_Identifier agent. This collected information has not undergone rigorous manual checking, and as a result, it may contain irrelevant or inaccurate details. We addressed this issue through Aspect_Analyst and Aspect_Reviewer agents.
@@
Liu et al. (Liu et al., 2024a) identified that the main drawback of role-based collaboration in multi-agent design involves increased communication overhead due to collaboration between agents, and the possibility of varying prices for agent services with different roles. In our pilot experiments, we found that making the interaction process of agents overly complex did not greatly improve the quality of the generated DR. Therefore, we referenced the single-path plan generator design pattern (Liu et al., 2024a), making each agent have only one subsequent step after completing its task, which is to submit the necessary information from the task execution results to the next agent.

### v3

@@ 3.3. Dataset Construction
As shown in Fig. 2, we constructed a dataset from SO posts, GitHub issues and discussions to facilitate and evaluate the DR generation for practical architecture decisions. To avoid potential data leakage issues, we constructed a new dataset, as the existing datasets may contain data that was used to train the LLMs used in this study as we discussed in Section 3.2. Note that, in our dataset, all posts, issues and discussions were created after the cut-off dates of the five LLMs used in this study. The dataset used in this study has been provided at (dataset).
@@ 3.3.1. Data Collection
For data collection, we used a keyword-based method to gather data from SO and GitHub. Specifically, we selected the general terms “architect*” (including “architecture”, “architecting”, and “architectural”) to search posts, issues and discussions. Such selection was also used in a recent study to retrieve architecture related SO posts (musengamana2023characterizing), and it indicates that these terms are effective in capturing more relevant data compared to other general keywords like “design*”. For SO posts, we utilized an SQL query via Stack Exchange Data Explorer interface (stackexchange) to collect posts with at least one answer from SO. For GitHub issues and discussions, we collected issues marked as “closed” with the reason specified as “completed”, and discussions marked as “answered”. Note that, we did not limit our search solely to the tags of SO posts, GitHub issues and discussions, as using tags alone may omit some important data related to software architecture topics (barua2014developers). Therefore, we considered both the body and title of posts, issues and discussions during our data search.
Additionally, we limited the retrieval of the posts, issues and discussions to dates after the knowledge cut-off dates of all LLMs. As presented in Table 1, gemini-1.0-pro has the latest cut-off date among the five LLMs we selected, that is, April 2023. Hence, we restricted the data retrieved from SO and GitHub to content after May 2023. We conducted data collection on June 18, 2024, gathering data from SO and GitHub that contained “architect*” in the content body, tags, or title, starting from May 1, 2023. In total, 692 SO posts, 32,614 issues, and 2,546 discussions were retrieved.
@@ 3.3.2. Data Review and Labeling
  • The posts, issues, and discussions must explicitly include an “architecture-related problem” ($\bm{P}$), and the corresponding “architecture decision” ($\bm{D}$) and “design rationale” ($\bm{DR}$). The relationship among $\bm{P}$, $\bm{D}$, and $\bm{DR}$, as defined in Equation 1, serves as the basis for our data extraction, ensuring that the required information (i.e., D1–D3 presented in Table 3) could be retrieved from each data entry.
@@
The formal data labeling was conducted by the first author. Prior to this, the first and fourth authors randomly selected 25 SO posts, 25 GitHub issues, and 25 GitHub discussions for a pilot labeling process to minimize personal bias and ensure adherence to the data labeling criteria. The inter-rater reliability between the two authors was measured by the Cohen’s Kappa coefficient (jacob1960coefficient), yielding values of 0.834. This result indicates an almost perfect agreement between the two authors (landis1977theMO). The second and third authors were then involved to resolve discrepancies between the pilot results of the first and fourth authors and reach a consensus on the four criteria for data labeling. Subsequently, the first author conducted the formal data labeling according to the four criteria. For SO posts, the first author reviewed and labeled all 692 posts, and identified 50 architecture-related posts for our research. During the data labeling process, 83 posts were excluded based on Criterion 1, whereas 358 and 204 posts were excluded based on Criteria 3 and 4, respectively. Given the substantial number of collected GitHub issues and discussions, and to maintain a balanced data volume between SO and GitHub, the two data sources, we decided to select 25 issues and 25 discussions that satisfied the inclusion criteria. Consequently, the final dataset comprises 100 entries, evenly split between the two data sources: 50 from SO and 50 from GitHub. The 25 issues and 25 discussions were obtained after the first author manually reviewed and labeled 586 GitHub issues and 784 GitHub discussions, respectively. A total of 114 issues and 21 discussions were excluded based on Criterion 2, 272 issues and 498 discussions were excluded based on Criterion 3, and 175 issues and 240 discussions were excluded based on Criterion 4. To assess whether the reviewed GitHub issues and discussions are representative of the collected data, we conducted statistical estimation based on the review size and the total number of collected data entries (israel1992determining). As a result, the margin of error was calculated to be 4.01% for the GitHub issues and 2.91% for the GitHub discussions at the 95% confidence level. The data labeling results were reviewed multiple times by the first, second, third, and fourth authors. Discrepancies were addressed through discussions to reach a consensus. Specifically, the first author presented the reasons for the inclusion or exclusion of certain data. The four authors then analyzed and discussed the reasons based on the criteria for data labeling, ultimately reaching an agreement. The pilot and formal data labeling results are recorded in our dataset (dataset).
Table 2 presents the taxonomy of eight types of architecture decisions with examples from our dataset, developed through the constant comparison method (stol2016grounded). Within the 100 architecture decisions collected from our dataset, Implementation Decisions (25%) are observed with the highest frequency, followed by Technology Decisions (15%) and Integration Decisions (14%). Architecture Pattern Decisions (12%), Data Decisions (12%), and Component Decisions (11%) occur less frequently, whereas Design Pattern Decisions (6%) and Deployment Decisions (5%) are the least frequent.
Table 2. Taxonomy of architecture decisions
Type Example Count
Implementation Decision If you want to auto-register the dependencies, use a different library or implement it yourself - e.g. take the domain Assembly, enumerate all types that end with “Query” and register them as scoped (SO #77843622) 25
Technology Decision From what I get, SignalDB should work fine with them (MPA and the Islands architecture). (Discussion #626) 15
Integration Decision I suggest writing all API controllers and authentication logic in the server project. Create a models project and reference this in both the client and server projects (SO #77815059) 14
Architecture Pattern Decision You would have to go for another architectural style, e.g. Onion Architecture. (Discussion #112) 12
Data Decision Instead, we should support ref in foreign key constraint expression - both at the model and column level. (Issue #8062) 12
Component Decision The second option is to pull Path.State (and all child feature States) into a separate module. (Discussion #2792) 11
Design Pattern Decision In Dart, we can use "implements" on a concrete class, so it is now my general pattern to make a concrete class for the repository, and merely “implement” that for my mock repository for testing. (SO #77579165) 6
Deployment Decision Approach: maintain a map or comparable structure of prisma client instances in the nextjs app / node server instance in memory, one for each db / schema instance requested (Discussion #2090) 5
@@ 3.3.3. Data Extraction
According to Equation 1, we established a set of data items for data extraction, as presented in Table 3. Data items D1 and D2 are provided as part of the prompt for LLMs and LLM-based agents to generate DR. D3 serves as a reference for evaluation in results analysis, as further elaborated in Section 3.5. The data extraction criteria are as follows:
@@
The first and fourth authors randomly selected 5 SO posts, 5 GitHub issues, and 5 GitHub discussions to conducted a pilot data extraction. In case of any discrepancies, the second and third authors were involved in reaching a consensus. The results show that the three data items in Table 3 can be extracted from our dataset. Subsequently, the first author conducted the formal data extraction from the labeled dataset. Each extracted data item was reviewed multiple times by the first, second, third, and fourth authors to ensure accuracy. Any inconsistencies between the four authors were discussed according to the criteria of data extraction to achieve a consensus. The data extraction results are recorded in our dataset (dataset).
@@
D1 Architecture Problem ($\bm{P}$) The key point(s) of the architecture problem in SO posts, GitHub issues, or GitHub discussions If you wish to access Console, Router, or Controller via the internet, you must open ports to the public internet and assign a public DNS record. This opens Ziti services to DDoS.
D2 Architecture Decision ($\bm{D}$) The key point(s) of the architecture decision in SO posts, GitHub issues, or GitHub discussions I propose implementing UDP hole-punching on the Controller and Router and Tunneler so they can see each other without opening the Controller and Router ports to the Internet. And since the Tunneler can now see the Controller and Router, you can also generate JWT locally on the Tunneller/Client.
D3 Design Rationale ($\bm{DR}$) The key point(s) of design rationale that consists of the arguments for the architecture decision in SO posts, GitHub issues, or GitHub discussions UDP hole-punching in products such as ZeroTier and Nebula that may be useful for OpenZiti: Argument_1: The supernode is a separate stand-alone service that does not influence sensitive services, such as the controller/router that does the main work and handles and enforces zero-trust, routing, and handle management. So, if the supernode is hacked, the other components will not be compromised. Argument_2: There are many supernodes across the Internet, so it is hard to DDoS all of them simultaneously. Argument_3: Even if all supernodes are DDoSed and down, the peers that already established connection do not care about the supernode and continue working without an issue. Argument_4: Only one port is opened, and you must be concerned about being open on the public Internet to one service that is not viral and acceptable if not working for hours or days.
@@ 3.4. Design Rationale Generation
To answer RQ1–RQ3 (see Section 3.1), we designed prompts for LLMs and LLM-based agents to generate DR based on the data items defined in Table 3. The DR generated with three prompting strategies is recorded in the dataset (dataset), which are further discussed in Section 4.
@@ 3.4.2. CoT (RQ2)
CoT allows LLMs to progressively demonstrate their thought process before arriving at a final answer, enhancing their reasoning ability and the transparency of their decisions (jason2022chain). Therefore, in addition to “Architecture Problem” and “Architecture Decision”, we also provided LLMs with a framework for analyzing the rationale in prompts. We designed several versions of the framework for LLMs to analyze the rationale and conducted pilot experiments on 10 randomly selected samples from our dataset. According to the results of the pilot experiments, we identified the most efficient CoT framework with four steps shown in Fig. 4. In Step 1, we ask the LLMs to understand architecture decisions based on the context of architecture problems. In Step 2 and Step 3, we ask the LLMs to analyze the advantages and disadvantages of the architecture decisions, respectively. Finally, in Step 4, we ask the LLMs to present the DR based on the trade-off of the advantages and disadvantages.
It is important to note that we did not provide any other examples in prompts for LLMs. This is because the main purpose of providing additional examples in CoT is to help the model understand the required reasoning process (jason2022chain). However, the results of the pilot experiments demonstrate that LLMs can generate DR step by step according to the designed CoT framework without any examples, which was further confirmed in the formal experiments. In addition, introducing examples in prompts increases the length of input text, which raises the difficulty for LLMs to understand architecture problems and leads to their generated DR being potentially influenced by specific examples.
@@ 3.4.3. LLM-based Agents (RQ3)
To answer RQ3 in Section 3.1, we designed five LLM-based agents (i.e., Aspect_Identifier, Information_Collector, Aspect_Analyst, Aspect_Reviewer, and Trade-off_Analyst) by utilizing the MetaGPT framework (hong2023metagpt). MetaGPT utilizes an assembly line approach to assign different roles to various agents, effectively breaking down complex tasks into subtasks that require collaboration among multiple agents. According to the study of Liu et al. (liu2024agent), the design pattern of MetaGPT is role-based cooperation, which has been employed in many SE studies (e.g., (jin2024mare; widyasari2024codeagent)). Currently, no research has reached a widespread consensus on the role delineation for generating DR. Thus, we approached a process of generating DR from the perspective of architects (jansen2008documenting; arman2018recovering), breaking it down into a set of subtasks assigned to five distinct agents.
@@
  • Aspect_Identifier agent is responsible for identifying the most relevant and important aspects related to certain architecture decisions. Specifically, Aspect_Identifier agent receives the architecture problem and the corresponding architecture decision, and then provides up to six of the most relevant aspects, along with their definitions. We set this upper limit at six aspects according to the results of the pilot experiments on 10 samples from our dataset. The results indicated that without such a restriction, LLMs tend to enumerate too many aspects, which can make it difficult to maintain a clear focus in the subsequent analysis of DR. Additionally, this excessive enumeration increases the costs for LLM-based agents. As a result, we found that limiting the aspects to six is considered a reasonable compromise, ensuring both depth of analysis and a concentrated focus. Then, each defined aspect is transmitted to Information_Collector agent.
@@
  • Information_Collector agent is in charge of gathering relevant information through search engines. We initially intended to enhance the information retrieval capabilities of Information_Collector agent by using Retrieval-Augmented Generation (RAG) (gao2024retrieval) technique with a unified database. However, the diversity of architecture problems prevented us from providing a single database that could adequately address all issues. Consequently, we chose to dynamically retrieve relevant knowledge from search engines based on specific problems. Specifically, Information_Collector agent receives architecture problems, the corresponding architecture decisions, and aspects predefined by Aspect_Identifier agent. Based on each aspect, Information_Collector generates the three most relevant keywords for search engine queries. We chose DuckDuckGo as our search engine as it offers search results without personalized interference and provides a high level of information reliability (hannak2013measuring). For each keyword, Information_Collector agent retrieves the top 20 URLs returned by the search engine. Therefore, for each aspect, a total of 60 relevant URLs are collected (i.e., 3 $\times$ 20 URLs). To prevent the agent from using the original SO posts, GitHub issues, or GitHub discussions as reference information (i.e., data leakage issues), we discard any collected background knowledge whose URLs appear in the dataset and replace them with alternative sources. Then, the agent stores the webpage content from these URLs in Facebook AI Similarity Search (FAISS), a library designed for fast nearest-neighbor search and dense vector clustering (douze2024faiss). By using the similarity search function of FAISS, Information_Collector agent selects the 10 most relevant pieces of information from all the collected webpages that are most similar to the given architecture problems and architecture decisions. The 10 pieces of information are then transmitted to Aspect_Analyst agent as background knowledge along with the aspects and their definitions identified by Aspect_Identifier agent. This collected information has not undergone rigorous manual checking, and as a result, it may contain irrelevant or inaccurate details. We addressed this issue through Aspect_Analyst and Aspect_Reviewer agents.
@@
Liu et al. (liu2024agent) identified that the main drawback of role-based collaboration in multi-agent design involves increased communication overhead due to collaboration between agents, and the possibility of varying prices for agent services with different roles. In our pilot experiments, we found that making the interaction process of agents overly complex did not greatly improve the quality of the generated DR. Therefore, we referenced the single-path plan generator design pattern (liu2024agent), making each agent have only one subsequent step after completing its task, which is to submit the necessary information from the task execution results to the next agent.